### v3

@@ 9.3 Proof of Lemma 3
Let $\{\lambda_{1},\lambda_{2},\dots,\lambda_{N_{V}}\}$ be the set of distinct eigenvalues of $V$. By the spectral theorem, $V$ can be decomposed as $V=U\Lambda U^{\top}$, where the columns of $U$ consist of orthonormal eigenvectors of $V$, denoted by $\{u_{11},\dots,u_{1n_{1}},u_{21},\dots,u_{2n_{2}},\dots,u_{N_{V}1},\dots,u_{N_{V}n_{N_{V}}}\}$, where $n_{1},\dots,n_{N_{V}}$ are the algebraic multiplicity of the eigenvalues respectively. Since $V$ is a symmetric matrix, the eigenvectors for a basis of $\mathbb{R}^{d}$ and we have $\sum_{i=1}^{N_{V}}n_{i}=d$. We can write $x$ as a linear combination $x=\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}u_{ij}$, where $\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}^{2}=1$. Define $a_{i}\coloneqq\sum_{j=1}^{n_{i}}w_{ij}^{2}$, by direct computation, we have $\sum_{i=1}^{N_{V}}a_{i}=1$ and $\|x\|_{V}^{2}=\sum_{i=1}^{N_{V}}\lambda_{i}a_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{N_{V}}\lambda_{i}^{-1}a_{i}$
@@
$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{N_{V}}a_{i}=1,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{N_{V}}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$ (10)
To satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices. Hence, $a_{i}$ can only be non-zero for at most two distinct eigenvalues, denoted by $\lambda_{i}$ and $\lambda_{j}$, where $i<j$ and $\lambda_{i}>\lambda_{j}$. Namely, the solution to (10) lies in the subspace spanned by the eigenvectors corresponding to $\lambda_{i}$ and $\lambda_{j}$. Let $y=\|x\|_{V}^{2}$, we have $a_{i}\lambda_{i}+a_{j}\lambda_{j}=y$ and $a_{i}\lambda_{i}^{-1}+a_{j}\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$. Note that $\sum_{i=1}^{N_{V}}a_{i}=a_{i}+a_{j}=1$, by direct computation, the closed form of $\|x\|_{V^{-1}}^{2}$ is given by $\|x\|_{V^{-1}}^{2}=\frac{1}{\lambda_{i}}+\frac{1}{\lambda_{j}}-\frac{y}{\lambda_{i}\lambda_{j}}\eqqcolon f(y,\lambda_{i},\lambda_{j})$. Clearly, we have
@@ 10.1 Proof of Proposition 3
By Propostion 1, we have
${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right]$
@@
${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]$
@@ 10.2 Proof of Proposition 4
Since ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta^{\prime}),\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})=\iota_{t}\beta^{RLS}_{t}(\delta^{\prime}).$ (11)
In the following, we drop the dependence on $\delta^{\prime}$ for notation simplicity. We have
@@
$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}.$
Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}$.
The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle ineuqlity of the $V_{t}^{-1}$-norm, and the concentration condition (11). As a result, we have
$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS},$
$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t},$
$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
Combing all terms above, we have
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
@@ 10.3 Proof of Theorem 1
$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
@@
$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t}^{RLS}(\delta^{\prime})=\beta_{T}^{RLS}(\delta^{\prime})$, we obtain
$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t}^{RLS}(\delta^{\prime})\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T}^{RLS}(\delta^{\prime}).$
@@
10.4 Proof of Proposition 5

### v5

@@ 10.3 Proof of Lemma 3
Let $\{\lambda_{1},\lambda_{2},\dots,\lambda_{N_{V}}\}$ be the set of distinct eigenvalues of $V$. By the spectral theorem, $V$ can be decomposed as $V=U\Lambda U^{\top}$, where the columns of $U$ consist of orthonormal eigenvectors of $V$, denoted by $\{u_{11},\dots,u_{1n_{1}},u_{21},\dots,u_{2n_{2}},\dots,u_{N_{V}1},\dots,u_{N_{V}n_{N_{V}}}\}$, where $n_{1},\dots,n_{N_{V}}$ are the algebraic multiplicity of the eigenvalues respectively. Since $V$ is a symmetric matrix, the eigenvectors form a basis of $\mathbb{R}^{d}$ and we have $\sum_{i=1}^{N_{V}}n_{i}=d$. We can write $x$ as a linear combination $x=\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}u_{ij}$, where $\sum_{i=1}^{N_{V}}\sum_{j=1}^{n_{i}}w_{ij}^{2}=1$. Define $a_{i}\coloneqq\sum_{j=1}^{n_{i}}w_{ij}^{2}$, by direct computation, we have $\sum_{i=1}^{N_{V}}a_{i}=1$ and $\|x\|_{V}^{2}=\sum_{i=1}^{N_{V}}\lambda_{i}a_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{N_{V}}\lambda_{i}^{-1}a_{i}$
@@
$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\forall i,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{N_{V}}a_{i}=1,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{N_{V}}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{N_{V}}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$ (9)
To satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices. Hence, $a_{i}$ can only be non-zero for at most two distinct eigenvalues, denoted by $\lambda_{i}$ and $\lambda_{j}$, where $i<j$ and $\lambda_{i}>\lambda_{j}$. Namely, the solution to (LABEL:eq:KKT) lies in the subspace spanned by the eigenvectors corresponding to $\lambda_{i}$ and $\lambda_{j}$.
Let $y=\|x\|_{V}^{2}$, we have $a_{i}\lambda_{i}+a_{j}\lambda_{j}=y$ and $a_{i}\lambda_{i}^{-1}+a_{j}\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$. Note that $\sum_{i=1}^{N_{V}}a_{i}=a_{i}+a_{j}=1$, by direct computation, the closed form of $\|x\|_{V^{-1}}^{2}$ is given by $\|x\|_{V^{-1}}^{2}=\frac{1}{\lambda_{i}}+\frac{1}{\lambda_{j}}-\frac{y}{\lambda_{i}\lambda_{j}}\eqqcolon f(y,\lambda_{i},\lambda_{j})$.
Clearly, we have
@@ 11.1 Proof of Proposition 3
By Proposition 1, we have
${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\geq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\geq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\right]$
@@
${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\notin\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}|\mathcal{F}_{t}\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\notin\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}|\mathcal{F}_{t}\right]$
@@ 11.2 Proof of Proposition 4
Since ${\theta}^{\star}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT}$, it holds that
$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS},\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}=\iota_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$ (10)
We have
@@
$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$.
The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle inequality of the $V_{t}^{-1}$-norm, and the concentration condition (10). As a result, we have
$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS},$
$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}},$
$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}.$
Combining all terms above, we have
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}.$
@@ 11.3 Proof of Theorem 1
$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
@@
$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}=\beta_{T,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}$, we obtain
$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}.$
@@
12 Regret Decomposition of POFUL
To discuss the relationship between our method and those based on optimism (Abbasi-Yadkori et al., 2011; Agrawal and Goyal, 2013; Abeille et al., 2017), we decompose the regret of POFUL into three terms, and sketch how they are bounded separately.
Let $x_{t}^{\star}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$ be the optimal action and $\widetilde{x}_{t}$ be the action chosen by POFUL, the regret decomposes as
$\mathcal{R}(T)$ $=\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$
$=\underbrace{\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle}_{\mathcal{R}^{PE}(T)}+\underbrace{\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle}_{\mathcal{R}^{PVT}(T)}+\underbrace{\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle}_{\mathcal{R}^{RLS}(T)}\,.$
$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator. Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate. By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda_{\text{reg}})$. Hence, the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which, via the Cauchy-Schwarz inequality, is characterized by the radius of the confidence ellipsoid $\mathcal{E}^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$. $\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}$ rather than using $\widehat{\theta}_{t}$ as OFUL would. Similar to $\mathcal{R}^{RLS}(T)$, the upper bound for this term is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$. As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$. Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound. For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$. However, it is known that, see for example (Agrawal and Goyal, 2013; Abeille et al., 2017; Hamidi and Bayati, 2020b), such selection could be problematic for bounding $\mathcal{R}^{PE}(T)$.
$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term. In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution, which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$. In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$.
For OFUL, this term is bounded properly with high probability by the optimistic selection of OFUL actions [^1] The term is further split into two parts as shown in the first few lines of proof of Proposition 4 in Appendix 11.2, with one term bounded by the optimism and the other term controlled by the $V_{t}$ distance between $\theta^{\star}$ and $\widetilde{\theta}_{t}$.. For LinTS, the only known analysis due to Agrawal and Goyal (2013) and Abeille et al. (2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$. The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ to enforce exploration. For example, Abeille et al. (2017) demonstrate that by using inflation, LinTS samples optimistic $\widetilde{\theta}_{t}$ with a probability greater than a constant, which means that the inequality $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$ holds with a constant probability.
For Greedy, there is no general theoretical guarantee for bounding the pessimism term without imposing additional structural assumptions. Our approach to tackle this challenge distinguishes us from methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017). Interestingly, Abeille et al. (2017) conjectured that non-optimistic samples may provide beneficial exploration and help control the growth of regret.
By relaxing the requirement for optimistic samples, we avoid inflating the posterior, which in turn prevents a $\sqrt{d}$-gap in the regret. Instead, we introduce a measure of the “quality” of POFUL actions with respect to the regret and develop a computable upper bound for this quality using geometric information in the data. This approach yields a data-driven regret bound for POFUL that matches the minimax optimal regret in some settings. Furthermore, it allows for the construction of variants of LinTS and Greedy with a provable frequentist regret bound that is minimax optimal up to only a constant factor.
13 Discrete Action Space
This section introduces the method for establishing an upper bound on $\alpha_{t}$ in the context of discrete action spaces. It turns out that the process involves comparing confidence interval lengths across all potentially optimal actions.
$L_{t}(x)\coloneqq\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\leq\langle{x},{{\theta}^{\star}}\rangle\leq\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}\eqqcolon U_{t}(x).$
the set of potentially optimal actions is given by $C_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS})$. Similarly, the set of actions that might be chosen by POFUL is given by $\mathcal{C}_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT})$, where we let $\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\coloneqq\iota_{t}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$. The inflation parameter $\iota_{t}$ turns out to control the conservativeness when eliminating actions.
$\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS})}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{PVT})}\|x\|_{V_{t}^{-1}}}.$ (11)
14 Example Cases and empirical validations of Theorem 2
To better understand what Theorem 2 implies, we discuss some special cases in this section.
Case 1: a pure exploration regime.
When the decision-maker doesn’t care about the regret and adopts a pure exploration algorithm that plays actions in all directions sequentially, we expect $\lambda_{i}(V_{t})=\mathcal{O}(t)$ for all $i=1,\dots,d$. Then $\widehat{\alpha}_{t}\leq C$ for some constant $C>0$. Specifically, if $V_{t}=DI_{d}$ for some constant $D>0$, we have $\|x\|_{V_{t}^{-1}}=D^{-1}\|x\|=D^{-1}$ for all $x\in{\mathcal{S}}_{d-1}$ and hence $\widehat{\alpha}_{t}=1$
Case 2: linear structures.
When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$ (empirically validated below). To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small. The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding uncertainty ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1. This provides a plausible explanation for the empirical success of LinTS and Greedy.
Proposition 5
Suppose $\lim_{t\to\infty}\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}}{\|\widehat{\theta}_{t}\|^{2}}=\lambda_{1}(V_{t})$, it holds that $\lim_{t\to\infty}\widehat{\alpha}_{t}=1$.